What volume of oxygen, measured under room conditions, can be obtained from the thermal decomposition of 8.2g of calcium nitrate?
A- 600cm^3
B-1200cm^3
C-150cm^3
D-300cm^3
The solution needed with an explanation.
Thanks
1
Expert's answer
2020-07-15T06:42:13-0400

Thermal decomposition of calcium nitrate:

2 Ca(NO3)2 → 2 CaO + 4 NO2 + O2

M (Ca(NO3)2) = 164.09 g/mol

n = m/M

n = 8.2 / 164.09 = 0.05 mol

According to the equation, and the mole ratio:

n (O2) = 1/2 n (Ca(NO3)2) = 0.05 / 2 = 0.025 mol

The molar volume of any gas is 24 dm3 mol-1 at room conditions

0.025 x 24 = 0.600 dm3  or 600 cm3

Correct answer: A 600 cm3